OpenStudy (anonymous):

if u=1-2tan3x, what is du equal to?

OpenStudy (anonymous):

use chain rule

OpenStudy (anonymous):

d/dx (tan 3x)= sec^2(3x) d/dx(3x)= sec^2(3x)3